What is the volume of 7.50 mol krypton at STP?
1
Expert's answer
2012-01-10T10:48:39-0500
PV=nRT
V=(nRT)/P
According to STP:
R=8,31 (The gas
constant)
P=100 kPa
T= 273.15
K
V=(7,50*8,31*273,15)/100=170,24 L